OpenStudy (anonymous):

A winter jacket is marked down 75% from its original price. According to the new price tag, it now costs $52.50. What was the original price?

OpenStudy (anonymous):

$210 original price

OpenStudy (mathmale):

@purpledinosaurs: Remember: guide the other person towards finding his or her own answer. Emphatically do NOT just provide the answer without explanation or involvement of the other person.

OpenStudy (mathmale):

Let x = original price. If the markdown was 75%, then this person is paying just 25% of the original price. Write an equation, involving x (the original price), 0.25 (the fraction of the original price that this person is paying for the jacket), and the new price. Solve for x. Show your work, please.

OpenStudy (skullpatrol):

$$\frac{\text{percent of change}}{{100}}=\frac{\text{change in price}}{\text{original price}}$$

OpenStudy (mathmale):

Original price: 1.00 x Markdown: 75%, or 0.75 Price actually paid: (1.00-0.75)*x=$52.50 or: (0.25)x +$52.50 Please solve this for x. Express your answer in dollars ($).